# Università della Svizzera italiana, Switzerland

### v1

@@ 3 Design
The goal of this study is to conduct a differentiated replication of the work by Wei et al. (2023a) to investigate the impact of AQLM, the quantization technique summarized in Section 2.3, on the performance of LLMs when applied to the code generation task. As previously explained, AQLM allows the exploration of extreme quantization levels, down to 2 bits, which may unlock new possibilities in the deployment of LLM-based code generators on constrained hardware environments. The context of the study consists of two state-of-the-art code LLMs, i.e., CodeLlama (Roziere et al., 2023) and DeepSeek-Coder (Guo et al., 2024), and two multilingual code generation benchmarks, i.e., MultiPL-E (Cassano et al., 2023) and McEval (Chai et al., 2024).
In particular, the study aims at addressing the following research questions (RQs):
@@
RQ1: How does low-bit quantization affect the model’s code generation ability? In RQ1, we aim to investigate the effect of low-bit quantization on the model’s ability to generate code starting from a textual description. As compared to the work by Wei et al. (2023a), which mostly focuses on 8-bit quantization, we experiment with 8, 4, 3, and 2-bit quantization. The goal is to carefully study the trade-off between lost in code generation accuracy and saving in terms of memory footprint. The RQ1’s findings will provide insights into the extent to which LLM-based code generators can be compressed without significantly compromising their performance.
@@
RQ2: Which impact does the calibration dataset have on model performance? In RQ2, we analyze the impact of the calibration dataset on model performance. The calibration dataset is used by the quantization technique to learn how to best approximate the original model’s weights. We experiment with various types of calibration datasets, including code-specific ones, which may be more suitable in the context of code generation and help in minimizing the loss in performance. Such an analysis is particularly important for the quantization of models targeting software engineering tasks and has not been conducted in the work of Wei et al. (2023a) where the authors focus solely on examining the extent to which the size of the calibration dataset, built on CodeSearchNet (Husain et al., 2019), affects quantization performance.
@@
3.1 Code-Specific Large Language Models Used in the Study
We focus on two state-of-the-art code LLMs, namely CodeLlama and DeepSeek-Coder. Both these LLMs have shown their effectiveness when applied to code generation tasks (Li et al., 2023a; Coignion et al., 2024; Ren et al., 2024).
3.1.1 CodeLlama (Roziere et al., 2023)
CodeLlama is a family of open-source LLMs tailored for coding applications, built on the general-purpose Llama 2 (Touvron et al., 2023) LLM. CodeLlama features Llama 2 models further trained on a corpus of 500 billion tokens, including code and natural language. Multiple versions of these models are publicly accessible111https://huggingface.co/codellama and are available in various sizes, with parameters ranging from 7B to 70B. Additionally, CodeLlama is available in three specialized formats: A general-purpose version for coding, an Instruct version optimized for instruction tuning, and a Python-specialized variant. CodeLlama has been selected as a representative model of code-related LLMs given its former successful applications in the automation of various code-related tasks (Weyssow et al., 2023; Huang et al., 2024; Sun et al., 2024a). In our study, we exploit the general-purpose CodeLlama versions featuring 7B, 13B, and 34B parameters. While we attempted to experiment with the 70B version, the computational cost would have been too high for our infrastructure
3.1.2 DeepSeek-Coder (Guo et al., 2024)
@@ 3.1 Code-Specific Large Language Models Used in the Study
3.2 Quantization Technique
As previously explained, we rely on the state-of-the-art method for quantization proposed by Egiazarian et al. (2024), namely AQLM. To clarify the specifics of our study, let us briefly review the fundamental workings of AQLM. Given a codebook consisting of candidate vectors and M being the number of available codebooks, AQLM encodes groups of g consecutive weights from each model layer by expressing them as the sum of vectors selected from M distinct codebooks. A model can be quantized at different levels of precision, measured in bits per parameter, based on the values assigned to g (number of weights to quantize together), M (number of codebooks), and B (number of bits per codebook). The choice of these values may have an impact on quantization time and, above all, on the final memory footprint and accuracy of the resulting quantized model.
The specific configurations we use for each experimented quantization level are reported in Table 1. A few clarifications are needed about these configurations. First, the values for the g, M, and B parameters are derived from the targeted quantization level. For example, setting g = 8, M = 4, B = 15 results in 8 bits per model’s parameters, while lowering M = 2 halves the precision to 4 bits. Second, given a parameters’ configuration, the obtained precision slightly varies among the models. For example, CodeLlama 7B quantized using g=8, M=1, and B=15, results in an average of 2.02 bits per parameter, compared to 1.92 bits per parameter for the 33B version. Finally, for the smallest model (i.e., DeepSeek-Coder 1B), which is only used in the context of RQ3 with the extreme 2-bit quantization, we had to lower B to 14 (rather than 15) to obtain a quantization level close to 2 bits (i.e., 2.05 bits per parameter).
@@ 3.2 Quantization Technique
Besides the specific AQLM parameters’ configuration, we had to create the calibration dataset used by the technique to minimize the quantization error (see Section 2.3).
@@
We create three different calibration datasets since they will be the focus of RQ2. For each of them, we selected 1,024 samples from RedPajama, each having a sequence length of 4,096 tokens, mimicking what was done by Egiazarian et al. (2024) when presenting AQLM. This means that from each of the 1,024 documents selected from RedPajama (e.g., a Wikipedia article), we only extract 4,096 consecutive tokens. The three variations of the calibration dataset differ in the “nature” of the documents they feature:
Random dataset: We randomly sampled the 1,024 samples from all the documents in the RedPajama dataset. Such a dataset simulates the choice of a “general purpose” calibration dataset, which may or may not be optimal in the context of code generation. Table 2 shows the distribution of the selected samples for both families of models (i.e., CodeLlama and DeepSeek-Coder). Note that we had to build different calibration datasets for the two families of LLMs since they exploit different tokenizers. Thus, a document may feature tokens for CodeLlama and for DeepSeek-Coder, with .
@@
Code dataset: In this case, the whole 1,024 instances come from the GitHub subset of RedPajama, with 512 being Java files and 512 being Python files.
| Source | Count |
| Common Crawl | 834 |
| C4 | 76 |
| Github | 62 |
| Arxiv | 28 |
| Wikipedia | 20 |
| StackExchange | 4 |
| Source | Count |
| Common Crawl | 835 |
| C4 | 70 |
| Github | 61 |
| Arxiv | 30 |
| Wikipedia | 24 |
| StackExchange | 4 |
@@ 3.3 Evaluation Dataset
Similarly to previous studies (Chen et al., 2021; Wei et al., 2023a; Cassano et al., 2024), we evaluate the model’s code generation ability by providing a specification as input and expecting a valid implementation as output. More specifically, the code generation task considers a function’s signature and documentation (i.e., natural language description) as a specification and its body as the expected implementation.
@@
We rely on two multilingual benchmarks: MultiPL-E (Cassano et al., 2023) and McEval (Chai et al., 2024). The first includes two popular code generation datasets, i.e., HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). Among these two datasets, we only used HumanEval, which was found to be more challenging than MBPP (Cassano et al., 2023). HumanEval originally included 164 hand-written Python prompts along with canonical solutions and unit test cases for evaluating model generations. The MultiPL-E benchmark retains 161 of these Python tasks, 158 of which have also been translated into Java. Note that three tasks were not translated because they used a Pythonic syntax not directly mappable into Java constructs. McEval is a collection of human-annotated coding tasks spanning different programming languages. Each problem has a signature, a docstring with an exhaustive description of the program requirements, a list of test cases, and a label indicating the level of difficulty (easy, middle, or hard) of the code to implement. We extracted the Python and Java problems from this benchmark, which amounted to 50 and 53 samples, respectively. We removed 8 problems from the Python samples that overlap with the HumanEval dataset, thus reducing the total number of Python samples to 42. The final Python (Java) problems feature 23 (30) easy tasks, 10 (13) middle-level samples, and 9 (10) hard problems. Our decision of focusing on Python and Java is due to their popularity and to the fact that they are considered high-resource languages, namely languages for which LLMs can exploit massive training material, thus usually exhibiting good code generation performances. This cannot really be said for low-resource languages (e.g., R, Racket), which would have introduced an additional challenge for the LLMs besides the extreme quantization we experiment with.
3.4 Study Procedure and Data Analysis
@@
To answer RQ1 we start by quantizing CodeLlama 7B and DeepSeek-Coder 7B via AQLM (Egiazarian et al., 2024) by leveraging the “Random” calibration dataset. The quantization is performed at 8, 4, 3, and 2 bits following the configurations in Table 1. The resulting models and the baseline (i.e., the non-quantized model) are then tested against the Java and Python samples from the two benchmarks to see how quantization affects model performance. In particular, given a code generation task, the candidate code (i.e., the one generated by an LLM) is validated against task-specific unit tests and labeled as “pass” if it satisfies all assertions or “fail” otherwise. As done by Wei et al. (2023a), as a performance metric, we use pass@1, which assesses the ability of the model to predict a correct solution with a single attempt (Chen et al., 2021) (i.e., the model is only allowed to output a single candidate solution). To account for the nondeterminism of the LLM, we invoke the model 20 times per code generation problem using a temperature of 0.2 and a maximum output length of 1.024 tokens for MultiPL-E problems and 1.500 tokens for the McEval benchmark, ensuring an output window capable of handling the more demanding requirements of McEval. As anticipated, at each of the 20 invocations, the model is only allowed to generate a single candidate solution. Finally, we compute the pass@1 score for each code generation task, i.e., the percentage of cases (out of 20) in which the LLM produced a correct solution with the given single attempt. Finally, we report the average score across all code generation tasks.
@@
To answer RQ2, we quantized CodeLlama 7B and DeepSeek-Coder 7B again at the same quantization levels used in RQ1 (i.e., 8, 4, 3, and 2 bits) using, however, the two additional calibration datasets presented in Section 3.2 (i.e., “Mixed” and “Code” dataset). Given the goal of RQ2 (i.e., investigating the impact on the performance of using different calibration datasets), in this context the baseline is represented by the quantized models used in RQ1, which exploited the “Random” calibration dataset.
Thus, we compare the pass@1 score achieved by the models quantized with the “Mixed” and “Code” calibration datasets against one of the models quantized with the “Random” dataset. We exploit the same statistical analysis previously described for RQ1 (i.e., adjusted -values resulting from McNemar’s test, and OR as effect size).
@@
Finally, in RQ1 and RQ3, we also experiment with an additional recently proposed strategy to further boost the performance of quantized models (Tseng et al., 2024; Egiazarian et al., 2024; Leconte et al., 2024). Such a strategy, known as “end-to-end fine-tuning”, can potentially extend the boundaries of the Pareto optimality, particularly when quantizing at low bit-size precision, such as 2 bits. The basic idea behind this additional fine-tuning (details in Appendix A of Egiazarian et al. (2024)) is to use the non-quantized LLM as a teacher model from which its quantized version can distill knowledge in order to minimize differences in their predictions on a given dataset. We run the end-to-end fine-tuning on the calibration dataset employed for the quantization process and re-used the same hyperparameter configuration by Egiazarian et al. (2024): 5 training epochs, a learning rate of 1e-5 combined with the Adam optimizer (Zhang, 2018), and an early stopping mechanism to prevent overfitting. We discuss the effect that this further “optimization strategy” has on the performance of quantized models at 3 and 2-bit levels. We do not present this analysis in RQ2 since our only goal there is to isolate the impact of the type of calibration dataset.
4 Study Results
| Symbol | p-value |
In this section, we report and discuss the results of our study to address the research questions formulated in Section 3. When reporting the results of the statistical tests, we adopt the notation in Table 3 for the significance level.
4.1 How does low-bit quantization affect the model’s code generation ability?
Fig. 1 provides a handy graphical representation of the memory saving (-axis) versus the relative performance loss in terms of pass@1 score (-axis) when comparing models quantized at different bit precisions (8, 4, 3, and 2 bits per parameter) and the baseline models using fp16 precision. The blue lines represent CodeLlama 7B while the orange lines DeepSeek-Coder 7B. The left charts relate to the Python benchmarks (from MultiPL-E and McEval) while the right one to the Java benchmarks.
@@ 4.1 How does low-bit quantization affect the model’s code generation ability?
Table 4 reports the detailed numbers on top of which Fig. 1 has been built: For each LLM subject of RQ1 (i.e., CodeLlama and DeepSeek-Coder 7B) we show, both for the baseline fp16 precision model as well as for all its quantized versions, (i) their memory footprint in terms of GB, (ii) the pass@1 score they achieved on both the Python and the Java benchmarks; and (iii) the results of the statistical tests (i.e., adjusted -value and OR), in which we compare each quantized model against the baseline. For example, looking at CodeLlama 7B, the fp16-precision baseline requires 13.48GB of memory to achieve a 29.8% of pass@1 score on the Python MultiPL-E benchmark, while its 4-bit quantized version only uses 4.00GB of memory with a pass@1 of 29.1%. This difference in performance is not statistically significant. When mapping this back to Fig. 1 (a) (blue line in the left chart), we can indeed see that the 4-bit quantized model, when compared to the fp16-precision baseline, allows achieving a 70% of memory reduction (i.e., ) at the rather small cost of a relative -2% (i.e., ) decrease in pass@1 score.
Looking at Fig. 1, the first conclusion that can be made is that, independently from the model and programming language, 4-bit quantization is a safe bet, ensuring high memory reduction with very similar performance in terms of code generation. The result of the 4-bit CodeLlama model on the Java McEval benchmark is the only exception to this finding, with a statistically significant decrease of pass@1 by 4%. To determine whether this behavior is due to the higher complexity of the benchmark, we checked the difficulty level of each prompt for which the base model returned a correct implementation while the quantized model did not. We discovered that 34 failed generations refer to easy problems, 16 to middle-level tasks, and 14 to hard tasks, which is consistent with the distribution of problem levels in the benchmark (30 easy, 10 middle, and 12 hard prompts). After a manual inspection, we found that the quantized model mainly failed due to assertion errors or calls to non-existent APIs.
@@
As anticipated in Section 3.4 we also experimented with end-to-end fine-tuning to boost the performances at 3 and 2-bit quantization, as suggested by (Egiazarian et al., 2024). Table 5 presents a comparison of the pass@1 score achieved by the 3- and 2-bit precision models, both with (dark grey) and without (light grey) end-to-end fine-tuning. For reference, the fp16-precision baseline is reported as well. In this case, the results of the statistical tests refer to the comparison between a quantized model with end-to-end fine-tuning versus the same quantized models without end-to-end fine-tuning. In other words, a statistically significant -value indicates a significant boost of pass@1 score given by the additional fine-tuning.
@@
4.2 Which impact does the calibration dataset have on model performance?
@@
For 8-bit and 4-bit models, we do not observe a statistically significant difference in pass@1 score between models quantized with different calibration datasets. This holds both for CodeLlama and DeepSeek-Coder and in both languages (Python and Java). This finding reveals that code models are robust to the calibration dataset provided at quantization time for target precisions greater or equal to 4 bits per parameter. The only exceptions are the CodeLlama 4-bit model on the McEval Java benchmark and the DeepSeek-Coder 4-bit model on the McEval Python benchmark, which are compared with a “Random” quantization that already performs better than the fp16-precision baseline. The 4-bit CodeLlama model, instead, shows a statistically significant improvement on the McEval Python benchmark after the “Mixed” calibration, thus recovering from the performance gap with the baseline observed in Section 4.1. On the other hand, we notice that 3-bit and 2-bit precision models are more sensible to the samples provided in the calibration dataset. Indeed, both models show a statistically significant improvement in performance when calibration datasets feature code samples, suggesting their need to better approximate the model weights at extreme quantization levels. Thanks to code-related calibration datasets, a 3-bit quantization might be considered an option in very resource-constrained hardware devices. However, when comparing the 3-bit models quantized using the “Mixed” dataset against the fp16-precision baseline, the difference in pass@1 score is still statistically significant.
Similarly, the 2-bit models continue to suffer from a significant performance gap despite major improvements brought by the code-related calibration datasets. This calls for additional research on optimizing calibration datasets for code-related tasks, which could help enhance performance even further. Indeed, we only experimented with three variants of the calibration datasets, which, due to the unavailability of the training sets used for CodeLlama and DeepSeek-Coder, we cannot guarantee to be representative of the training instances (a condition that is expected to help the quantization process (Egiazarian et al., 2024)).
Although our results do not show a clear winner between the two datasets, providing both code and technical language generally yields better performance for both models. This is particularly evident in the CodeLlama family of models, as they are more susceptible to sophisticated natural language requirements, such as those found in the McEval dataset. We believe including code and domain-specific natural language in the calibration dataset can help models that frequently struggle to follow complex instructions, like CodeLlama.
4.3 How does extreme quantization affect model accuracy across different model sizes?
@@
For what concerns the fine-tuned variants, we observe that training after quantization can considerably help smaller models to achieve better accuracy. For example, when fine-tuning DeepSeek-Coder 1B, the performance decrease on Java (MultiPL-E) passes from -77.1% to -49%, while the one on Python goes from -51.1% to -23.6%. In general, we observe that fine-tuning helps reduce the quantization error on smaller models. However, as the number of model parameters increases, we notice that the impact of fine-tuning gradually diminishes and eventually becomes counterproductive.
5 Implications of our Findings

### v2

@@ 3. Design
The goal of this study is to conduct a differentiated replication of the work by Wei et al. (Wei et al., 2023) to investigate the impact of AQLM on the performance of LLMs when applied to the code generation task. Also, we aim at studying the impact of different types of calibration datasets on the quantization process, an aspect not covered in the original study (Wei et al., 2023). As previously explained, AQLM allows the exploration of extreme quantization levels, down to 2 bits, which may unlock new possibilities in the deployment of LLM-based code generators on constrained hardware environments. The context of the study consists of two state-of-the-art code LLMs, i.e., CodeLlama (Roziere et al., 2023) and DeepSeek-Coder (Guo et al., 2024), and two multilingual code generation benchmarks, i.e., MultiPL-E (Cassano et al., 2023) and McEval (Chai et al., ).
We aim at addressing the following research questions (RQs):
@@
RQ1: How does low-bit quantization affect the model’s code generation ability? In RQ1, we aim to investigate the effect of low-bit quantization on the model’s ability to generate code starting from a textual description. As compared to the work by (Wei et al., 2023), which mostly focuses on 8-bit quantization, we experiment with 8, 4, 3, and 2-bit quantization. The goal is to carefully study the trade-off between loss in code generation accuracy and savings in terms of memory footprint. The RQ1’s findings will provide insights into the extent to which LLM-based code generators can be compressed without significantly compromising their performance.
@@
RQ2: What impact does the calibration dataset have on model performance? In RQ2, we analyze the impact of the calibration dataset on model performance. The calibration dataset is used by the quantization technique to learn how to best approximate the original model’s weights. We experiment with various types of calibration datasets, including code-specific ones, which may be more suitable in the context of code generation and help in minimizing the loss in performance. Such an analysis is particularly important for the quantization of models targeting software engineering tasks and has not been conducted in the work of Wei et al. (Wei et al., 2023) where the authors focus solely on examining the extent to which the size of the calibration dataset, built on CodeSearchNet (Husain et al., 2019), affects quantization performance.
@@
3.1. Code-Specific LLMs Used in the Study
We focus on two state-of-the-art code LLMs, namely CodeLlama and DeepSeek-Coder. Both these LLMs have shown their effectiveness when applied to code generation tasks (Li et al., 2023; Coignion et al., 2024; Ren et al., 2024).
3.1.1. CodeLlama (Roziere et al., 2023)
CodeLlama is a family of open-source LLMs tailored for coding applications, built on the general-purpose Llama 2 (Touvron et al., 2023) LLM. CodeLlama features Llama 2 models further trained on a corpus of 500 billion tokens, including code and natural language. Multiple versions of these models are publicly accessible and are available in various sizes, with parameters ranging from 7B to 70B. Additionally, CodeLlama is available in three specialized formats: A general-purpose version for coding, an Instruct version optimized for instruction tuning, and a Python-specialized variant. CodeLlama has been selected as a representative model of code-related LLMs given its former successful applications in the automation of various code-related tasks (Weyssow et al., 2023; Huang et al., 2024; Sun et al., 2024a). In our study, we exploit the general-purpose CodeLlama versions featuring 7B, 13B, and 34B parameters. While we attempted to experiment with the 70B version, the computational cost would have been too high for our infrastructure. However, given the results obtained with the 34B version, we believe that similar trends would be observed with the 70B version as well.
3.1.2. DeepSeek-Coder (Guo et al., 2024)
@@ 3.1. Code-Specific LLMs Used in the Study
3.2. Quantization Technique
As previously explained, we rely on the state-of-the-art method for quantization proposed by Egiazarian et al. (Egiazarian et al., 2024), namely AQLM. To clarify the specifics of our study, let us briefly review the fundamental workings of AQLM. Given a codebook consisting of candidate vectors and M being the number of available codebooks, AQLM encodes groups of g consecutive weights from each model layer by expressing them as the sum of vectors selected from M distinct codebooks.
A model can be quantized at different levels of precision, measured in bits per parameter, based on the values assigned to g (number of weights to quantize together), M (number of codebooks), and B (number of bits per codebook). The choice of these values may have an impact on quantization time, inference latency, and on the final memory footprint and accuracy of the quantized model. In this work we focus on memory footprint and accuracy. Inference latency is a different optimization objective, which we leave as future work. The specific configurations we use for each experimented quantization level are reported in Table 1. A few clarifications are needed about these configurations. First, the values for the g, M, and B parameters are derived from the targeted quantization level. For example, setting g = 8, M = 4, B = 15 results in 8 bits per model’s parameters, while lowering M = 2 halves the precision to 4 bits. Second, given a parameters’ configuration, the obtained precision slightly varies among the models. For example, CodeLlama 7B quantized using g=8, M=1, and B=15, results in an average of 2.02 bits per parameter, compared to 1.92 bits per parameter for the 33B version. Finally, for the smallest model (i.e., DeepSeek-Coder 1B), which is only used in the context of RQ3 with the extreme 2-bit quantization, we had to lower B to 14 (rather than 15) to obtain a quantization level close to 2 bits (i.e., 2.05 bits per parameter).
@@ 3.2. Quantization Technique
Besides the specific AQLM parameters’ configuration, we had to create the calibration dataset used by the technique to minimize the quantization error (see Section 2).
@@
We create three different calibration datasets since they will be the focus of RQ2. For each of them, we selected 1,024 samples from RedPajama, each having a sequence length of 4,096 tokens, mimicking what was done by Egiazarian et al. (Egiazarian et al., 2024) when presenting AQLM. This means that from each of the 1,024 documents selected from RedPajama (e.g., a Wikipedia article), we only extract 4,096 consecutive tokens. The three variations of the calibration dataset differ in the “nature” of the documents they feature:
Random dataset: We randomly sampled the 1,024 samples from all the documents in the RedPajama dataset. Such a dataset simulates the choice of a “general purpose” calibration dataset, which may or may not be optimal in the context of code generation. Table 2 shows the distribution of the selected samples for both families of models. Note that we had to build different calibration datasets for the two families of LLMs since they exploit different tokenizers. Thus, a document may feature tokens for CodeLlama and for DeepSeek-Coder, with .
@@
Code dataset: The whole 1,024 instances come from the GitHub subset of RedPajama, with 512 being Java files and 512 Python files.
| Source | CodeLlama | DeepSeek-Coder |
| Common Crawl | 834 | 835 |
| C4 | 76 | 70 |
| Github | 62 | 61 |
| Arxiv | 28 | 30 |
| Wikipedia | 20 | 24 |
| StackExchange | 4 | 4 |
@@ 3.3. Evaluation Dataset
Similarly to previous studies (Chen et al., 2021; Wei et al., 2023; Cassano et al., 2024), we evaluate the model’s code generation ability by providing a specification as input and expecting a valid implementation as output. More specifically, the code generation task considers a function’s signature and documentation (i.e., natural language description) as a specification and its body as the expected implementation.
@@
We rely on two multilingual benchmarks: MultiPL-E (Cassano et al., 2023) and McEval (Chai et al., ). The first includes two popular code generation datasets, i.e., HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). Among these two datasets, we only used HumanEval, which was found to be more challenging than MBPP (Cassano et al., 2023). HumanEval originally included 164 hand-written Python prompts along with canonical solutions and unit test cases for evaluating model generations. The MultiPL-E benchmark retains 161 of these Python tasks, 158 of which have also been translated into Java. Note that three tasks were not translated because they used a Pythonic syntax not directly mappable into Java constructs. McEval is a collection of human-annotated coding tasks spanning different programming languages. Each problem has a signature, a docstring with an exhaustive description of the program requirements, a list of test cases, and a label indicating the level of difficulty (easy, middle, or hard) of the code to implement. We extracted the Python and Java problems from this benchmark, which amounted to 50 and 53 samples, respectively. We removed 8 problems from the Python samples that overlap with the HumanEval dataset, thus reducing the total number of Python samples to 42. The final Python (Java) problems feature 23 (30) easy tasks, 10 (13) middle-level samples, and 9 (10) hard problems. Our decision of focusing on Python and Java is due to their popularity and to the fact that they are considered high-resource languages, namely languages for which LLMs can exploit massive training material, thus usually exhibiting good code generation performances. This cannot really be said for low-resource languages (e.g., R, Racket), which would have introduced an additional challenge for the LLMs besides the extreme quantization we experiment with.
3.4. Study Procedure and Data Analysis
@@
To answer RQ1 we start by quantizing CodeLlama 7B and DeepSeek-Coder 7B via AQLM (Egiazarian et al., 2024) by leveraging the “Random” calibration dataset. The quantization is performed at 8, 4, 3, and 2 bits following the configurations in Table 1. Note that we experiment with only the 7B versions to contain the computational cost of the study while still being able to observe the impact of quantization on reasonably sized models. The resulting models and the baseline (i.e., the non-quantized model) are then tested against the Java and Python samples from the two benchmarks to see how quantization affects model performance. In particular, given a code generation task, the candidate code (i.e., the one generated by an LLM) is validated against task-specific unit tests and labeled as “pass” if it satisfies all assertions or “fail” otherwise. As done by Wei et al. (Wei et al., 2023), as a performance metric, we use pass@1, which assesses the ability of the model to predict a correct solution with a single attempt (Chen et al., 2021) (i.e., the model is only allowed to output a single candidate solution). To account for the nondeterminism of the LLM, we invoke the model 20 times per code generation problem using a temperature of 0.2 and a maximum output length of 1,024 tokens for MultiPL-E problems and 1,500 tokens for the McEval benchmark, ensuring an output window capable of handling the more demanding requirements of McEval. As anticipated, at each of the 20 invocations, the model is only allowed to generate a single candidate solution. Finally, we compute the pass@1 score for each code generation task, i.e., the percentage of cases (out of 20) in which the LLM produced a correct solution with the given single attempt. Finally, we report the average score across all code generation tasks.
@@
To answer RQ2, we quantized CodeLlama 7B and DeepSeek-Coder 7B again at the same quantization levels used in RQ1 (i.e., 8, 4, 3, and 2 bits) using, however, the two additional calibration datasets presented in Section 3.2 (i.e., “Mixed” and “Code” dataset). Given the goal of RQ2 (i.e., investigating the impact on the performance of different calibration datasets), the baseline is represented by the quantized models used in RQ1, which exploited the “Random” calibration dataset. Thus, we compare the pass@1 achieved by the models quantized with the “Mixed” and “Code” calibration datasets against one of the models quantized with the “Random” dataset. We use the same statistical tests described for RQ1.
@@
Finally, in RQ1 and RQ3, we also experiment with an additional recently proposed strategy to further boost the performance of quantized models (Tseng et al., 2024; Egiazarian et al., 2024; Leconte et al., 2024). Such a strategy, known as “end-to-end fine-tuning”, can potentially extend the boundaries of the Pareto optimality, particularly when quantizing at low bit-size precision, such as 2 bits. The basic idea behind this additional fine-tuning (details in Appendix A of (Egiazarian et al., 2024)) is to use the non-quantized LLM as a teacher model from which its quantized version can distill knowledge in order to minimize differences in their predictions on a given dataset. We run the end-to-end fine-tuning on the calibration dataset employed for the quantization process and re-used the same hyperparameter configuration by (Egiazarian et al., 2024): 5 training epochs, a learning rate of 1e-5 combined with the Adam optimizer (Zhang, 2018), and an early stopping mechanism to prevent overfitting. We discuss the effect that this further “optimization strategy” has on the performance of quantized models at 3 and 2-bit levels. We do not present this analysis in RQ2 since our only goal there is to isolate the impact of the type of calibration dataset.
4. Study Results
In this section, we report and discuss the results of our study to address the research questions formulated in Section 3. When reporting statistical tests, we adopt the following notation for the significance level: • indicates a -value being ();
* (); ** (); and *** ().
How does low-bit quantization affect the model’s code generation ability?
@@ 4. Study Results
Table 3 reports for each LLM subject of RQ1 (i.e., CodeLlama and DeepSeek-Coder 7B) and for both the baseline fp16 precision model as well as for all its quantized versions: (i) their memory footprint in terms of GB, (ii) the pass@1 score they achieved on both the Python and the Java benchmarks; and (iii) the results of the statistical tests (i.e., adjusted -value and OR), in which we compare each quantized model against the baseline. For example, looking at CodeLlama 7B, the fp16-precision baseline requires 13.48GB of memory to achieve a 29.8% of pass@1 score on the Python MultiPL-E benchmark, while its 4-bit quantized version only uses 4.00GB of memory with a pass@1 of 29.1%. This difference in performance is not statistically significant. The 4-bit quantized model, when compared to the fp16-precision baseline, allows achieving a 70% of memory reduction (i.e., ) at the rather small cost of a relative -2% (i.e., ) decrease in pass@1 score.
Looking at Table 3, the first conclusion that can be made is that, independently from the model and programming language, 4-bit quantization is a safe bet, ensuring high memory reduction with very similar performance in terms of code generation. The result of the 4-bit CodeLlama model on the Java McEval benchmark is the only exception to this finding, with a statistically significant decrease of pass@1 by 4%. To determine whether this behavior is due to the higher complexity of the benchmark, we checked the difficulty level of each prompt for which the base model returned a correct implementation while the quantized model did not. We discovered that 34 failed generations refer to easy problems, 16 to middle-level tasks, and 14 to hard tasks, which is consistent with the distribution of problem levels in the benchmark (30 easy, 10 middle, and 12 hard prompts). After a manual inspection, we found that the quantized model mainly failed due to assertion errors or calls to non-existent APIs.
@@
As anticipated in Section 3.4 we also experimented with end-to-end fine-tuning to boost the performances at 3 and 2-bit quantization, as suggested by Egiazarian et al. (Egiazarian et al., 2024). Table 4 presents a comparison of the pass@1 score achieved by the 3- and 2-bit precision models, both with (dark grey) and without (light grey) end-to-end fine-tuning. For reference, the fp16-precision baseline is reported as well. In this case, the results of the statistical tests refer to the comparison between a quantized model with end-to-end fine-tuning versus the same quantized models without end-to-end fine-tuning. In other words, a statistically significant -value indicates a significant boost of pass@1 score given by the additional fine-tuning.
@@
Which impact does the calibration dataset have on model performance?
@@
For 8-bit and 4-bit models, we do not observe a statistically significant difference in pass@1 score between models quantized with different calibration datasets. This holds both for CodeLlama and DeepSeek-Coder and in both languages (Python and Java). This finding reveals that code models are robust to the calibration dataset provided at quantization time for target precisions greater or equal to 4 bits per parameter. The only exceptions are the CodeLlama 4-bit model on the McEval Java benchmark and the DeepSeek-Coder 4-bit model on the McEval Python benchmark, which are compared with a “Random” quantization that already performs better than the fp16-precision baseline. The 4-bit CodeLlama model, instead, shows a statistically significant improvement on the McEval Python benchmark after the “Mixed” calibration, thus recovering from the performance gap with the baseline observed in Section 4. On the other hand, we notice that 3-bit and 2-bit precision models are more sensible to the samples provided in the calibration dataset. Indeed, both models show a statistically significant improvement in performance when calibration datasets feature code samples, suggesting their need to better approximate the model weights at extreme quantization levels. Thanks to code-related calibration datasets, a 3-bit quantization might be considered an option in very resource-constrained hardware devices. However, when comparing the 3-bit models quantized using the “Mixed” dataset against the fp16-precision baseline, the difference in pass@1 score is still statistically significant.
Similarly, the 2-bit models continue to suffer from a significant performance gap despite major improvements brought by the code-related calibration datasets. This calls for additional research on optimizing calibration datasets for code-related tasks, which could help enhance performance even further. Indeed, we only experimented with three variants of the calibration datasets, which, due to the unavailability of the training sets used for CodeLlama and DeepSeek-Coder, we cannot guarantee to be representative of the training instances (a condition that is expected to help the quantization process (Egiazarian et al., 2024)). Although our results do not show a clear winner between the two datasets, providing both code and technical language generally yields better performance for both models. This is particularly evident in the CodeLlama family of models, as they are more susceptible to sophisticated natural language requirements, such as those found in the McEval dataset. We believe including code and domain-specific natural language in the calibration dataset can help models that frequently struggle to follow complex instructions, like CodeLlama.
How does extreme quantization affect model accuracy across different model sizes?
@@
For what concerns the fine-tuned variants, we observe that training after quantization can considerably help smaller models to achieve better accuracy. For example, when fine-tuning DeepSeek-Coder 1B, the performance decrease on Java (MultiPL-E) passes from -77.1% to -49%, while the one on Python goes from -51.1% to -23.6%. We observe that fine-tuning helps reduce the quantization error on smaller models. However, as the number of model parameters increases, we notice that the impact of fine-tuning gradually diminishes and eventually becomes counterproductive.
5. Implications of our Findings